OpenStudy (anonymous):

The graph of a line passes through the two points (-2, 1) and (2, 1). What is the equation of the line written in general form? A: y - 1 = 0 B: x - y + 1 = 0 C: x + y - 1= 0

OpenStudy (tkhunny):

Have you considered either building the equation with the two-point form or perhaps you could substitute the points and see which ones work? Show your work!

OpenStudy (anonymous):

I have, I figured out that it does equal zero but the equations they show are confusing

OpenStudy (anonymous):

do we seriously have to promise medal and fan this is ridiculous

OpenStudy (tkhunny):

No, the chats are a wasteland. No one there cares about helping anyone. Sorry about that. Just post your question AND your work and you should get help quickly in most cases. How are the equation confusing? A: y - 1 = 0 (-2, 1) and (2, 1) Try y = 1 A: 1 - 1 = 0 -- That looks good.

OpenStudy (anonymous):

oh, I messed up I thought it was 0-1=1, I just messed everything up for myself. Thank you xD